OpenStudy (anonymous):

Does the equation represent a direct variation? if so identify the constant of variation. 5x=-3y

OpenStudy (anonymous):

The equation y/x = 6 states that y "varies directly as" x since the ratio of y to x (also written y:x) never changes. The number 6 in the equation y/x = 6 is called the constant of variation. The equation y/x = 6 can also be written in the equivalent form, y = 6x. That form shows you that y is always 6 times as much as x. For example y = x/3 the constant of variation is 1/3. Your equation does represent direct variation, so first rearrange to x= -3y/5, so whats the constant knowing the info above?

OpenStudy (anonymous):

@KCpaige are you following me?

OpenStudy (anonymous):

Sorry the equation i posted above is to solve for x, heres to solve for y. Well to solve for y heres the steps, -3y = 5x y = (-5/3) *x the constant of variation is -5/3 Do you understand? :)

OpenStudy (anonymous):

@KCpaige ??

OpenStudy (anonymous):

give me one second im reviewing it( was folding clothes)

OpenStudy (anonymous):

y = (-5/3)x or y = -5x/3 --> k = -5/3